The data from 200 machined parts are summarized as follows Depth of Bore Edge Condition Coarse Moderate Smooth Above Target Below Target 10 23 84 15 46 (a) What is the probability that à part selected has a moderate edge condition and a below-target bore depth? Round your answer to two decimal places (e.g. 98.76) (b) What is the probability that a part selected has a moderate edge condition or a below-target bore depth? Round your answer to two decimal places (e.g. 98.76). () What is the probability that a part selected does not have a moderate edge condition or des no havea below-tar (c) What is t bore depth? Round your answer to two decimal places (e.g. 98.76).
Answer #1
condition above below total
coarse 15 10 25
moderate 22 23 45
smooth 46 84 130
total 83 117 200

a)

P =23/200=0.115 ~0.12

b)

P=(45+117-23)/200=0.695~0.70

c)

P=1-23/200=0.88 ( please try 0.89 if this cormes wrong and revert)